OpenStudy (anonymous):

-4+5a+3b/6a - 5a +5b/3a

OpenStudy (anonymous):

like this? (-4+5a+3b)/6a - (5a +5b)/3a

OpenStudy (anonymous):

its -4 + 5a+ 3b/ 6a - 5a +5b / 3a I have to write as a simple fraction

OpenStudy (anonymous):

like this? \[\frac{ -4 + 5a+ 3b}{ \frac{ 6a - 5a +5b}{ 3a}}\]

OpenStudy (anonymous):

no the -4 is separate and then its-4 + 5a +3b 5a +5b __________ - ________ 6a 3a

OpenStudy (anonymous):

like this then? \[-4 + \frac{ 5a+ 3b}{ 6a } - \frac{ 5a +5b }{ 3a }\]

OpenStudy (anonymous):

to simplify we will need to turn the two fractions into 1 fraction, can you do that by using a common denominator?

OpenStudy (anonymous):

yes it is just like that. the common denominator that I came up with was 3 but only for one side. I am so lost

OpenStudy (anonymous):

no worries! I'll show ya

OpenStudy (anonymous):

so when working with fractions and finding a common denominator, I want you to take the term that I circle and multiply it by the term that I draw the arrow to: